OpenStudy (mokeira):

A farmer harvested 900 bags of maize. She sold 0.7 bags ad gave 0.1 of the remainder to charitable organization. She then kept the rest. How many bags of maize were kept?

OpenStudy (mokeira):

my answer=243bags

OpenStudy (triciaal):

total = 900 sold = 0.7 * 900 = 630 charitable = 0.1* (900 - 630) = 27 left = 0.9(270) = 243 agree

OpenStudy (mokeira):

awesome!